OpenStudy (anonymous):

I am off for today guys good night and here is 1 sum find area of surface obtained by y=2sqrt(x) 4<=x<=9 about x axis

OpenStudy (anonymous):

no1 knows?

OpenStudy (anonymous):

\[\huge \frac{ 76 }{ 3 }\]

OpenStudy (anonymous):

what equation did you use?

OpenStudy (anonymous):

what did you integrate

OpenStudy (anonymous):

what is the answer? given

OpenStudy (anonymous):

let me know how you solved ? O.o no answer is given

OpenStudy (anonymous):

iam wrong !

OpenStudy (anonymous):

w8 let me solve it

OpenStudy (anonymous):



OpenStudy (anonymous):

total area = sum of perimeters all the small tiny circles from x=4 to x=9

OpenStudy (anonymous):

so integrating from 4 to 9 in terms of y=

OpenStudy (anonymous):

\[\int\limits_{4}^{9}2\sqrt{x}\]

OpenStudy (anonymous):

and a correction for thickness.... times, \[\sqrt{1+y'(x)}dx\]

OpenStudy (anonymous):

the element has a thickness dx so, gotta account for that

OpenStudy (anonymous):

wait I didn't get that

OpenStudy (anonymous):

\[ S=2\pi\int_xy\sqrt{1+y'^2}dx \]

OpenStudy (anonymous):

that's the final equation?

OpenStudy (anonymous):